OpenStudy (anonymous):

I don't understand why the limit of x/sqrt(x^2 - 1) when x -> infinity is 1, and when x -> - infinity is -1 anybody ?

OpenStudy (phi):

First, as \( x \rightarrow -\infty\), we know the numerator x is negative and the denominator \( \sqrt{x^2 -1}\) is positive (we take the principal i.e. positive root) thus the limit must be negative. on the other hand, for \( x \rightarrow +\infty\), the limit will be positive.

OpenStudy (phi):

to find the numerical value of the limit, we divide top and bottom by x to get \[ \frac{1}{\frac{\sqrt{x^2-1}}{x}} \] for \(x \rightarrow +\infty\) we can replace x with \( \sqrt{x^2}\) (again we assume the principal root) and we have \[ \frac{1}{\sqrt{\frac{x^2-1}{x^2}}}= \frac{1}{\sqrt{1-\frac{1}{x^2}}}\] and as x approaches infinity 1/x^2 goes to zero, and we have \[ \lim_{x \rightarrow \infty} \frac{1}{\sqrt{1-\frac{1}{x^2}}}= \frac{1}{\sqrt{1}}= 1\] for \(x \rightarrow -\infty\) , we do the same thing, but we must write \[x = - \sqrt{x^2} \] to get the correct sign. \[ \lim_{x \rightarrow -\infty} \frac{1}{\frac{\sqrt{x^2-1}}{-\sqrt{x^2}}} \\ = -\frac{1}{\sqrt{1-\frac{1}{x^2}}}= -1 \]